OpenStudy (anonymous):

what is the solutionset of 3x+ y =10 if each variable represents a whole number ?

jhonyy9 (jhonyy9):

y=10-3x - so for x=1 --- y=10-3=7 (1,7)

jhonyy9 (jhonyy9):

and continue

OpenStudy (anonymous):

Thanks